# Gate-tunable nonlocal Josephson effect through magnetic van der Waals bilayers

### v1

@@ IV.3 Josephson current through F interlayers
Figure 7: Critical Josephson current as a function of the magnetic bilayer length $L$ for the case of the bilayer composed of a ferromagnetic and a normal layers with $h_{1}=h$ and $h_{2}=0$. Different curves correspond to different values of $\Delta\mu$. $h_{1}=0.1t_{x}^{S}$, other parameters are the same as at Fig.5.
The case of AF bilayer considered in the previous section is special in the sense that at $\Delta\mu=0$ there are no oscillations of the critical current due to $h_{1}=-h_{2}$. Then the sensitivity of the system to gating is the strongest. If the condition $h_{1}=-h_{2}$ is violated the critical current begins to oscillate as a function of the interlayer length. As an example we consider an interlayer composed of a ferromagnetic and a normal layers with $h_{2}=0$. The critical current as a function of the interlayer length is plotted in Fig. 7 for different values of $\Delta\mu$. It is seen that the critical current oscillates even at $\Delta\mu=0$. The oscillation period is in good agreement with analytical results, which at $t_{z}^{F}\ll(\Delta\mu,h)$ can be expanded as $\lambda_{1,2}\approx-h/v_{F}\mp h\Delta\mu/2v_{F}t_{z}^{F}$ with $\hat{w}_{1,2}\approx(1,\pm 1,\pm 1,1)^{T}$, which provide the leading contribution to the Josephson current, and $\lambda_{3,4}\approx(\mp 2t_{z}^{F}-h)/v_{F}\mp(h^{2}+\Delta\mu^{2})/4v_{F}t_{z}^{F}$ with $\hat{w}_{3,4}\approx(-1,\mp 1,\pm 1,1)^{T}$, which only make a minor contribution to the critical current.
Figure 8: Bilayer composed of a ferromagnetic and a normal layers. Critical Josephson current as a function of the difference between on-site energies of magnetic layers $\Delta\mu$. $L=100a$, other parameters are the same as at Fig.7.
In the considered case for small $\Delta\mu$ the dependence of the oscillation period on $\Delta\mu$ is only a correction to the main term, which is associated to $\lambda_{1,2}\approx-h/v_{F}$. At large $\Delta\mu\gg t_{z}^{F}$ the difference between on-site energies effectively works as an additional barrier between the layers (mismatch effect). Then it is obvious that in this case the Josephson current is just a sum of independent contributions of two individual layers, one of which is oscillating due to the presence of $h_{1}=h$ and the other one is non-oscillating due to $h_{2}=0$. Thus, even in this case one can control $0-\pi$ transitions by applying the gating voltage, see Fig. 8. In general, the Josephson current is sensitive to gating for a wide range of magnetic bilayers except for the case of a ferromagnetic bilayer with $h_{1}=h_{2}$ because in this case the oscillation periods of the pairs, which provide the leading contribution to the Josephson current, do not depend on $\Delta\mu$, as it is seen from the left column of Fig. 3.
@@ V Conclusions
In this paper a theory of proximity effect in ballistic bilayer vdW magnetic systems is developed. Possible types of proximity-induced Cooper pairs are studies and their spatial behavior is described analytically. The considered proximity effect in the vdW bilayer systems is shown to be uniquely characterized by the presence of non-local Cooper pairs, which spatial behavior in the non-superconducting regions is qualitatively different from the well-known proximity effect in conventional non-magnetic and magnetic non-superconducting materials. In particular, the oscillation period of spatial oscillations of such pairs in non-superconducting regions is shown to be sensitive to the difference between on-site energies of the monolayer composing the bilayer and, thus, can be controlled by applying a gating potential to one of the monolayers. This opens the perspective of implementing the gate-controlled $0-\pi$ transitions in Josephson junctions through weak links made of such bilayers.
@@ Acknowledgements.
G.A.B., D.S.R. and I.V.B. acknowledge the support from Theoretical Physics and Mathematics Advancement Foundation “BASIS” via the project No. 23-1-1-51-1. The analytical calculations in the framework of quasiclassical approach were supported by the Russian Science Foundation via the RSF project No. 24-12-00152. The numerical calculations of the Josephson current were supported by Grant from the ministry of science and higher education of the Russian Federation No. 075-15-2024-632.
Appendix A Derivation of the two-layer quasiclassical Eilenberger equation
Here we describe the main steps of the derivation of the quasiclassical Eilenberger equation used in the main text for studying penetration of local and nonlocal pairs into the non-superconducting bilayer. We begin with the derivation of the Gor’kov equation for the full quantum mechanical Green’s function. To do this let us consider a general bilayer system described by the following tight-binding Hamiltonian:
Here $\hat{c}_{\bm{i},\sigma}=(c_{\bm{i},\sigma}^{1},c_{\bm{i},\sigma}^{2})^{T}$ is a vector composed of annihilation operators for electrons belonging to the first and the second layers at site $\bm{i}$ in the plane of each layer and for spin $\sigma=\uparrow,\downarrow$. $\hat{\mu}=\left(\begin{matrix}\mu_{1}&0\\ 0&\mu_{2}\\ \end{matrix}\right)$ is the matrix composed of on-site energies of the first and the second layers, respectively. $t$ is the hopping element in plane of each layer and $t_{12}$ is the hopping element between the first and the second layers. $\bm{\hat{h}}=\left(\begin{matrix}\bm{h}_{1}&0\\ 0&\bm{h}_{2}\\ \end{matrix}\right)$ is the matrix containing the exchange field of both layers and $\hat{\Delta}_{\bm{i}}=\left(\begin{matrix}\Delta_{1,\bm{i}}&0\\ 0&\Delta_{2,\bm{i}}\\ \end{matrix}\right)$ is the matrix composed of superconducting order parameters of the layers, which can be zero if we consider a non-superconducting bilayer.
The Matsubara Green’s function in the two-layer formalism is a $8\times 8$ matrix in the direct product of spin, particle-hole and layer spaces. Introducing the two-layer Nambu spinor $\check{\psi}_{\bm{i}}=(\hat{c}_{{\bm{i}},\uparrow}^{1},\hat{c}_{\bm{i},\downarrow}^{1},\hat{c}_{\bm{i},\uparrow}^{2},\hat{c}_{\bm{i},\downarrow}^{2},\hat{c}_{\bm{i},\uparrow}^{1\dagger},\hat{c}_{\bm{i},\downarrow}^{1\dagger},\hat{c}_{\bm{i},\uparrow}^{2\dagger},\hat{c}_{\bm{i},\downarrow}^{2\dagger})^{T}$ we define the Green’s function as follows:
where $\langle T_{\tau}...\rangle$ means imaginary time-ordered thermal averaging. Introducing Pauli matrices in spin, particle-hole and layer spaces: $\sigma_{k}$, $\tau_{k}$ and $\rho_{k}$ ($k=0,x,y,z$) and operator $\hat{j}$ as
$\hat{j}\hat{c}_{\bm{i}}=\sum\limits_{<\bm{i}\bm{j}>}t\hat{c}_{\bm{j}}$ (20)
and expanding $\check{G}_{\bm{i}\bm{j}}(\tau_{1}-\tau_{2})=T\sum\limits_{\omega_{m}}e^{-i\omega_{m}(\tau_{1}-\tau_{2})}\check{G}_{\bm{i}\bm{j}}(\omega_{m})$ over Matsubara frequencies $\omega_{m}=\pi T(2m+1)$, one can obtain the Gor’kov equation for Green’s function $\check{G}_{\bm{i}\bm{j}}(\omega_{m})$. The derivation is similar to that described in Ref. 63. The resulting Gor’kov equation takes the form:
$G_{\bm{i}}^{-1}\check{G}_{\bm{i}\bm{j}}(\omega_{m})=\delta_{\bm{i}\bm{j}},$ (21)
$G_{\bm{i}}^{-1}=\tau_{z}\left(\hat{j}+\hat{\mu}-\check{\Delta}_{\bm{i}}i\sigma_{y}-\bm{\hat{h}}\check{\bm{\sigma}}\delta_{i_{x},0}+t_{12}\rho_{x}\right)+i\omega_{m}.$ (22)
where $\bm{R}=(\bm{i}+\bm{j})/2$ and the integration is over $\bm{r}=\bm{i}-\bm{j}$. To simplify further calculations and to present the Gor’kov equation in a more common form we define the following transformed Green’s function:
where and below subscript $\tau$ means that the explicit matrix structure corresponds to the particle-hole space. In the quasiclassical approximation we disregard terms of the order of $1/(\xi_{S}p_{F})\ll 1$ with respect to unity, where $p_{F}$ is the Fermi momentum, $\xi_{S}=v_{F}/|\Delta|$ is the superconducting coherence length and $\hbar=1$. For this reason we can approximate $\hat{\Delta}_{\bm{i}}\approx\hat{\Delta}(\bm{R})$. The neglected terms $\sim(\bm{\nabla}\hat{\Delta}_{\bm{i}})\bm{r}\sim|\Delta_{1,2}|/(p_{F}\xi_{S})$. In general, the same is valid for the exchange field if it is inhomogeneous in plane of the layer. With these approximations we obtain:
$\check{\tilde{G}}(\bm{R},\bm{p})=1$ (29)
Subtracting Eqs. (29) and (30) we come to the following equation:
$F([\hat{j},\check{\hat{G}}])=0$ (31)
Now we introduce quasiclassical $\xi$-integrated Green’s function:
[…]
where $\xi=-2t[\cos(p_{x}a)+\cos(p_{y}a)]$ and $a$ is the lattice constant. After the $\xi$-integration of Eq. (31) the resulting Eilenberger equation for the quasiclassical Green’s function takes the form:
$+i\bm{v}_{F}\bm{\nabla}\check{g}(\bm{R},\bm{p}_{F})=0,$ (33)
$\check{g}^{2}(\bm{R},\bm{p}_{F})=1.$ (34)

### v2

@@ IV.3 Josephson current through F interlayers
Figure 7: Critical Josephson current as a function of the magnetic bilayer length $L$ for the case of the bilayer composed of a ferromagnetic and a normal layers with $h_{1}=h$ and $h_{2}=0$. Different curves correspond to different values of $\Delta\mu$. $h_{1}=0.1t_{x}^{S}$, other parameters are the same as in Fig.5.
The case of the AF bilayer considered in the previous section is special in the sense that at $\Delta\mu=0$ there are no oscillations of the critical current due to $h_{1}=-h_{2}$. Then the sensitivity of the system to gating is the strongest. If the condition $h_{1}=-h_{2}$ is violated the critical current begins to oscillate as a function of the interlayer length. As an example we consider an interlayer composed of a ferromagnetic and normal layers with $h_{2}=0$. The critical current as a function of the interlayer length is plotted in Fig. 7 for different values of $\Delta\mu$. It is seen that the critical current oscillates even at $\Delta\mu=0$. The oscillation period is in good agreement with analytical results, which at $t_{z}^{F}\ll(\Delta\mu,h)$ can be expanded as $\lambda_{1,2}\approx-h/v_{F}\mp h\Delta\mu/2v_{F}t_{z}^{F}$ with $\hat{w}_{1,2}\approx(1,\pm 1,\pm 1,1)^{T}$, which provide the leading contribution to the Josephson current, and $\lambda_{3,4}\approx(\mp 2t_{z}^{F}-h)/v_{F}\mp(h^{2}+\Delta\mu^{2})/4v_{F}t_{z}^{F}$ with $\hat{w}_{3,4}\approx(-1,\mp 1,\pm 1,1)^{T}$, which only make a minor contribution to the critical current.
Figure 8: Bilayer composed of a ferromagnetic and a normal layers. Critical Josephson current as a function of the difference between on-site energies of magnetic layers $\Delta\mu$. $L=100a$, other parameters are the same as in Fig.7.
In the considered case for small $\Delta\mu$ the dependence of the oscillation period on $\Delta\mu$ is only a correction to the main term, which is associated to $\lambda_{1,2}\approx-h/v_{F}$. At large $\Delta\mu\gg t_{z}^{F}$ the difference between on-site energies effectively works as an additional barrier between the layers (mismatch effect). Then it is obvious that in this case the Josephson current is just a sum of independent contributions of two individual layers, one of which is oscillating due to the presence of $h_{1}=h$ and the other one is non-oscillating due to $h_{2}=0$. Thus, even in this case one can control $0-\pi$ transitions by applying the gating voltage, see Fig. 8. In general, the Josephson current is sensitive to gating for a wide range of magnetic bilayers except for the case of a ferromagnetic bilayer with $h_{1}=h_{2}$ because in this case the oscillation periods of the pairs, which provide the leading contribution to the Josephson current, do not depend on $\Delta\mu$, as seen from the left column of Fig. 3.
@@ V Conclusions
In this paper a theory of proximity effect in ballistic bilayer vdW magnetic systems is developed. Possible types of proximity-induced Cooper pairs are studies and their spatial behavior is described analytically. The considered proximity effect in the vdW bilayer systems is shown to be uniquely characterized by the presence of non-local Cooper pairs, whose spatial behavior in the non-superconducting regions is qualitatively different from the well-known proximity effect in conventional non-magnetic and magnetic non-superconducting materials. In particular, the oscillation period of spatial oscillations of such pairs in non-superconducting regions is shown to be sensitive to the difference between on-site energies of the monolayer composing the bilayer and, thus, can be controlled by applying a gating potential to one of the monolayers. This opens the perspective of implementing the gate-controlled $0-\pi$ transitions in Josephson junctions through weak links made of such bilayers.
@@ Acknowledgements.
G.A.B., D.S.R. and I.V.B. acknowledge the support from Theoretical Physics and Mathematics Advancement Foundation “BASIS” via the project No. 23-1-1-51-1. The analytical calculations in the framework of the quasiclassical approach were supported by the Russian Science Foundation via the RSF project No. 24-12-00152. The numerical calculations of the Josephson current were supported by Grant from the ministry of science and higher education of the Russian Federation No. 075-15-2024-632.